# Progressive Multi-Source Domain Adaptation for Personalized Facial Expression Recognition

### v1

@@ III-B Source Selection
Before progressive adaptation of the model $\Theta$ to the target domain, we aim to select the most suitable domains from a multi-source domain $\mathcal{S}$ to align them with the target domain $\mathbf{T}$, based on a curriculum-based approach. Initially, we prioritize the source domains with high transferability to align them with the target domain. This will help select source domains with feature distributions similar to the target domain. After aligning the feature distributions of these source domains, a source selection [^1] For algorithm, see Section I-A of the supplementary material will prioritize the next round of source domains for alignment. As adaptation (training) continues, the model $\Theta$ gradually learn to focus on various aspects of the feature distribution to improve transferability. Our approach involves learning a curriculum to prioritize different source domains. We hypothesize that source domains closest to the target domain in the feature space are the ones that are easier for the model to adapt to. We compute the cosine similarity between every source domain and the target domain in a mini-batch [58], where the domains are represented by their feature embeddings as:
@@
where .$\|.\|_{2}$ is the $\ell^{2}$-norm of the feature embeddings.
@@
where $\mathbf{M}^{\text{c}}(.)$ is the cosine similarity between every feature embedding in $\mathcal{S}$ and $\mathbf{T}$, $D$ represents the total number of sources domains, and $P$ is a dictionary (list) that stores all pair-wise distances indexed by the respective source domain information. To determine the stopping criteria for including source domains in the adaptation process, we apply a normalization procedure on similarity measures in $P$ to scale them to the range $[0,1]$ and we obtain $\widetilde{P}$, which is the scaled version of $P$. Subsequently, we apply a predetermined threshold $\gamma$ to $\widetilde{P}$ to limit the number of source domains. We formulate the process of selecting source domains as:
@@
where $\widetilde{\mathcal{S}}$ denote the subset of source domains that meet the criterion $(\widetilde{P}_{j}>\gamma)$ and as consequence, are selected to updating the model $\Theta$. For each subset $\mathcal{\widetilde{S}}$ satisfying this condition, we compute a supervised loss as:
@@ III-C Progressive Domain Adaptation of Target Subject
Domain Alignment. We further mitigate the divergence between the domains using Maximum mean discrepancy (MMD) [60], which estimates the disparity among two distributions in RKHS space. In our problem, we have three subject domains $(\mathcal{\widetilde{S}}_{a},\mathbf{\widehat{T}},\mathbf{R})$, we jointly calculate the pairwise distances between $(\mathcal{\widetilde{S}}_{a},\mathbf{\widehat{T}})$ and $(\mathcal{\widetilde{S}}_{a},\mathbf{R})$. For every new source, the disparity is calculated between the target domain, source domain, and replay domain which makes sure to eliminate the domain shift among them.
@@
Source Combine
Source-only
Sub-based (UDA) [18]
0.62
0.73
0.61
0.64
0.65
0.73
0.55
0.59
0.51
0.54
0.71
0.75
0.70
0.76
0.52
0.53
0.54
0.51
0.55
0.58
0.59
0.63
Multi-Source
M3SDA [35]
CMSDA [20]
SImpAI [37]
Sub-based [18]
Sub-basedtop-k [18]
0.67
0.93
0.80
0.93
0.93
0.66
0.47
0.69
0.69
0.71
0.61
0.81
0.55
0.84
0.86
0.58
0.87
0.75
0.66
0.87
0.55
0.53
0.52
0.60
0.88
0.50
0.84
0.81
0.76
0.92
0.67
0.57
0.71
0.84
0.86
0.56
0.54
0.61
0.55
0.77
0.54
0.74
0.59
0.62
0.84
0.67
0.70
0.56
0.66
0.68
0.60
0.70
0.65
0.71
0.83
@@ IV Experimental methodology
FER datasets, such as RAF-DB [61], or AffectNet [62] in the context of MSDA setting, often contain mixed individuals, that do not represent subject variations required to developed a subject-based adaptation method. Nevertheless, we evaluate our progressive MSDA technique for expression recognition on two widely used pain recognition datasets, Biovid part(A) [63], and UNBC-Master [64]. These datasets provide subject information with balanced class distributions, following the same experimental protocol established in [18]. Furthermore, we study the impact of gradually incorporating relevant source samples versus entire source subjects. It will elucidate the optimal strategy for enhancing performance and generalization capabilities in subject-based adaptation.
@@ IV-A Datasets
BioVid Heat and Pain (PartA) [63] The dataset comprises 87 subjects recorded in a controlled environment, where each subject is categorized into one of five classes: ”no pain” and four escalating pain levels labeled PA1 through PA4, representing increasing pain intensity. Previous studies have indicated that the lower pain intensities, particularly in the initial stages, did not elicit noticeable facial activities. It is recommended that the focus be on the ”no pain” and highest pain intensity categories. Our experiments concentrate on two classes: ”no pain” and the highest pain level, PA4. Each subject contributes 20 videos per class, lasting 5.5 seconds each. Following the findings in [65], which noted that PA4 does not display significant facial activity in the first 2 seconds of the video, we exclude frames from the initial 2 seconds. This ensures that only the latter part of the sequence, where the subject’s response to pain is more pronounced, is analyzed.
@@
UNBC-McMaster Shoulder Pain [64] comprises 25 subjects and includes 200 video sequences. Pain intensity for each frame is assessed using the PSPI scale [66], which ranges from 0 to 15. Given the substantial imbalance across pain intensity levels, we adopt the quantization strategy used in [67], where pain intensities are grouped into five discrete levels: 0 (no pain), 1 (intensity 1), 2 (intensity 2), 3 (intensity 3), 4 (intensities 4-5), and 5 (intensities 6-15).
IV-B Implementation Detail
In all experiments, we employ the ResNet18 backbone [68], which consists of the encoder $F$ and the discriminative component $C$. To adapt $F$ for subject-based MSDA, we follow the same protocol as  [18] to remove the first ReLU, the MaxPool layers, and the final 2D adaptive average pooling layer. In our experiments, the backbone is shared across every domain, followed by the shared classifier. The images are resized to 100$\times$100 resolution, and the model is trained with stochastic gradient descent (SGD) with a batch size of 16 and a learning rate of $10^{-4}$. The closest source selection criterion is based on the threshold ($\tau$), which is set to 0.8, and the $top_{s}$ is set to $40$ subjects through empirical evaluation [^3] Details are provided in Section III-B of the supplementary material. . For generating target pseudo-labels we set a threshold based on $\theta=\theta_{0}-\delta\left\lfloor\frac{e}{N}\right\rfloor$, the initial value of $\theta_{0}$ is set to 0.91, that was updated after every $N=20$, with the reduction value $\delta$ is set to 0.01. For the ACPL technique to generate reliable target PLs, we follow the same setting as Zeeshan et al. [18] and use a horizontal flip as an augmented version of the image. For replay relevant samples, we set $top_{n}$ and $N_{r}$ to 2000 samples, which are updated after each newly added source subject.
IV-C Baseline Methods
To evaluate the performance of our method, we define subjects as source and target domains. The first experiment was conducted on the BioVid dataset, where 77 subjects were treated as sources and adapted to the remaining ten target subjects. The following experiment is on the UNBC-McMaster dataset, which includes 20 subjects in the source domain adapted to the remaining five subjects in the target domain. To evaluate the efficacy of our model, we further experimented with a cross-dataset with 20 UNBC-McMaster sources adapted to 10 BioVid target subjects. We follow the previous work of Zeeshan et al. [18] to define the MSDA standard for pain recognition.
Source-combined: We first define the lower-bound, which is the traditional approach of training a model by using all the sources and testing the target subject. This approach is also known as source-only, as it does not adapt to the target data. The second experiment combines all subjects as before and then adapts to a target subject as in standard UDA.
Multi-source DA: We treat each subject as a separate domain while adapting to the target. We evaluated our method with three standard and one subject-based MSDA approaches moment matching for multi-source domain adaptation (M3SDA) [35], implicit alignment (SImpAl) [37], contrastive multi-source domain adaptation (CMSDA), and subject-based domain adaptation [18]. M3SDA technique was the baseline method for MSDA classification tasks that reduced the discrepancy based on the moment-matching approach between domains. CMSDA and SImpAI methods are based on generating the target PLs. Subject-based DA is the STA technique used in multi-source domain adaptation for pain estimation.
V Results and Discussion
V-A Comparison with State-of-the-Art
The result on the BioVid dataset is shown in Table I. In the source-only method, we follow the lower bound approach without any form of domain adaptation, where a model is trained on training subjects (sources) and evaluated on unlabeled target subjects. In the source-combined (blending) setting, we compare the result with the subject-based UDA method [18]. We compare our technique with four state-of-the-art MSDA methods. SImpAl, CMSDA, M3SDA, and subject-based MSDA. Our method achieves higher performance for all target subjects with an average accuracy of 0.88 that exceeds the baseline by a large margin. The closest result was with a sub-basedtop-k with an average of 0.83 with a similar performance in Sub-2 with 0.86 accuracy. Almost every model performance is impressive in subject Sub-1. However, our method yields remarkable performance in achieving the accuracy of 0.99, which is the same as Oracle.
The result on UNBC-McMaster is presented in Table II. Our approach outperforms all methods on every target subject adaptation, including source-only and state-of-the-art MSDA methods, achieving an average accuracy of 0.88, which is a gain from the previous subject-based method 0.86. In particular, there is a significant improvement in Sub-1 performance with an increase of 0.6 compared to the subject-based approach. For Sub-3, we match the performance with the subject-based method with 0.94 accuracy.
@@ V-A Comparison with State-of-the-Art
SImpAI
CMSDA
Sub-based
Sub-basedtop-k
@@ V-B Cross-Dataset Evaluation
Figure 3: Analysis of different techniques to select the relevant samples for the replay domain. No Preserve, at each increment step we only use the new subject without considering previously visited samples; Preserve Random Samples, select fixed random samples from previous domains; Closest k-means, select samples based on k-means clusters; Closest DBSCAN (100/subject), incorporate 100 samples from each visited source subject using DBSCAN; Ours, density-based selection of pertinent samples.
@@
Figure 6: (a) We compare four techniques to gradually introduce source data; i) Random Source Samples, ii) Closest Source Samples, iii) Closest Source Subjects (all), and iv) Closest Source subjects (ours). Every point represents an average of all the target subjects in BioVid. On the x-axis, each increment is equivalent to approximately. 2k samples or a single subject. (b) Loss when introducing a new source domain, adapting to BioVid target subject-10. (c) Time complexity comparison for training; All Previous Subjects, CMSDA, SImpAI, and P-MSDA (ours).
V-C Impact of Progressive Adaptation on Source Domains
In this experiment, we further study the alternative approach to gradually learning by introducing the closest source samples instead of progressively introducing entire source domains. Fig. 6 (a) compares the effectiveness of selecting the closest source subjects with the closest source samples, random source samples, and the combination of all previously seen source subjects. Initially, we selected random source samples from all source domains and gradually introduced subsets of 2000 images, equivalent to the number of training images in a single subject from the BioVid dataset. The alternative approach is to select the closest samples from all source domains for each target based on cosine similarity. We gradually introduced these samples while retaining all previously seen samples. Our results indicate that this method improved performance compared to randomly selecting source samples. Nevertheless, the model performed even better when introducing the closest subjects. This improvement can be attributed to preserving subject-specific features that maintain the contextual information within the same subject, which helped the model learn more consistent representations relevant to each target. While training with all closest subjects and accumulating previously visited subjects initially outperformed the other approaches. However, as more subjects were included over time, there was less focus on newer subjects, resulting in minimal gains in adaptation performance. On the other hand, when only training with the closest source subjects and retaining the most relevant previous samples, performance improved significantly over time. This approach also required less memory and computational resources since it did not involve saving all previously seen data, leading to faster convergence.

### v2

@@ III-B Source Selection
Before progressive adaptation of the model $\Theta$ to the target domain, we aim to select the most suitable domains from a multi-source domain $\mathcal{S}$ to align them with the target domain $\mathbf{T}$, based on a curriculum-based approach. Initially, we prioritize the source domains with high transferability to align them with the target domain. This will help select source domains with feature distributions similar to the target domain. After aligning the feature distributions of these source domains, a source selection [^1] For algorithm, see Section I-A of the supplementary material will prioritize the next round of source domains for alignment. As adaptation (training) continues, the model $\Theta$ gradually learn to focus on various aspects of the feature distribution to improve transferability. Our approach involves learning a curriculum to prioritize different source domains. We hypothesize that source domains closest to the target domain in the feature space are the ones that are easier for the model to adapt to. We compute the cosine similarity between every source and the target domain in a mini-batch [57], domains are represented by their feature embeddings as:
@@
where $\|.\|_{2}$ is the $\ell^{2}$-norm of the feature embeddings.
@@
where $\mathbf{M}^{\text{c}}(.)$ is the cosine similarity between every feature embedding in $\mathcal{S}$ and $\mathbf{T}$, $D$ represents the total number of source domains, and $P$ is a dictionary (list) that stores all pair-wise distances indexed by the respective source domain information. To determine the stopping criteria for including source domains in the adaptation process, we apply a normalization procedure on similarity measures in $P$ to scale them to the range $[0,1]$ and we obtain $\widetilde{P}$, which is the scaled version of $P$. Subsequently, we apply a predetermined threshold $\gamma$ to $\widetilde{P}$ to limit the number of source domains. We formulate the process of selecting source domains as:
@@
where $\widetilde{\mathcal{S}}$ denotes the subset of source domains that meets the criterion $(\widetilde{P}_{j}>\gamma)$ and as a consequence, it is selected to update the model $\Theta$. For each subset $\mathcal{\widetilde{S}}$ satisfying this condition, we compute a supervised loss as:
@@ III-C Progressive Domain Adaptation of Target Subject
Pseudo-label for Target Domain. To calculate the pseudo-labels for the target domain, motivated by the augmented confident pseudo-label (ACPL) technique presented in [18], we calculate the target labels by generating an augmented version $\widehat{\mathbf{x}}^{t}$ of each target sample $\mathbf{x}^{t}$ using a model $\Theta$. We then estimate the probabilities as $\mathbf{p}^{\text{t}}=\mbox{softmax}(\mathbf{x}^{t})$ and $\widehat{\mathbf{p}}^{\text{t}}=\mbox{softmax}(\widehat{\mathbf{x}}^{t})$, taking the average of two probabilities $a^{t}=(\widehat{\mathbf{p}}^{\text{t}}+\mathbf{p}^{\text{t}})/2$. The selection criteria to assign the pseudo-label to the target sample is determined by the confidence threshold $\tau$ based on $\tau=\tau_{0}-\delta\left\lfloor\frac{e}{N}\right\rfloor$, where $e$ the current epoch, $N$ represents the epoch after which $\tau$ decreases, $\delta$ is the reduction value, and $\lfloor\cdot\rfloor$ is the floor function that rounded down to the nearest value. We assign the pseudo-labels to the target samples $\mathbf{\widehat{T}}=(\hat{\mathbf{x}}^{\text{t}},\hat{y}^{\text{t}})$ if $a^{t}$ is greater than $\tau$. At each training step, we estimate the target domain loss as:
@@
Source Combine Source-only 0.62 0.61 0.65 0.55 0.51 0.71 0.70 0.52 0.54 0.55 0.59
DANN [60] 0.73 0.57 0.69 0.73 0.68 0.79 0.55 0.61 0.60 0.55 0.65
CDAN [61] 0.64 0.50 0.68 0.71 0.51 0.69 0.61 0.63 0.67 0.47 0.61
Sub-based (UDA) [18] 0.73 0.64 0.73 0.59 0.54 0.75 0.76 0.53 0.51 0.58 0.63
Multi-Source M3SDA [35] 0.67 0.66 0.61 0.58 0.55 0.50 0.67 0.56 0.54 0.67 0.60
CMSDA [20] 0.93 0.47 0.81 0.87 0.53 0.84 0.57 0.54 0.74 0.70 0.70
SImpAI [37] 0.80 0.69 0.55 0.75 0.52 0.81 0.71 0.61 0.59 0.56 0.65
Sub-based [18] 0.93 0.69 0.84 0.66 0.60 0.76 0.84 0.55 0.62 0.66 0.71
Sub-basedtop-k [18] 0.93 0.71 0.86 0.87 0.88 0.92 0.86 0.77 0.84 0.68 0.83
@@ IV Experimental methodology
FER datasets such as RAF-DB [62] or AffectNet [63] are widely used but contain mixed individuals, making them less suited for subject-based adaptation where subject variability is crucial. To address this, we evaluate our progressive MSDA approach on two benchmark pain recognition datasets—BioVid Part A [64] and UNBC-McMaster [65]—and extend our study to in-the-wild datasets, Aff-Wild2 [66] and Behavioral Ambivalence/Hesitancy (BAH) [67]. While BioVid, UNBC, and BAH explicitly provide subject information, Aff-Wild2 is adapted to a subject-based setup by treating each video as a unique subject, following the protocol in [18]. Furthermore, we analyze the impact of progressively incorporating relevant source samples versus full source subjects, shedding light on strategies that best enhance performance and generalization in subject-based adaptation.
@@ IV-A Datasets
BioVid Heat and Pain (PartA) [64] The dataset comprises 87 subjects recorded in a controlled environment, where each subject is categorized into one of five classes: ”no pain” and four escalating pain levels labeled PA1 through PA4, representing increasing pain intensity. Previous studies have indicated that the lower pain intensities, particularly in the initial stages, did not elicit noticeable facial activities. It is recommended that the focus be on the ”no pain” and the highest pain intensity categories. Our experiments concentrate on two classes: ”no pain” and the highest pain level, PA4. Each subject contributes 20 videos per class, lasting 5.5 seconds each. Following the findings in [68], which noted that PA4 does not display significant facial activity in the first 2 seconds of the video, we exclude frames from the initial 2 seconds. This ensures that only the latter part of the sequence, where the subject’s response to pain is more pronounced, is analyzed.
@@
UNBC-McMaster Shoulder Pain [65] comprises 25 subjects and includes 200 video sequences. Pain intensity for each frame is assessed using the PSPI scale [69], which ranges from 0 to 15. Given the substantial imbalance across pain intensity levels, we adopt the quantization strategy used in [70], where pain intensities are grouped into five discrete levels: 0 (no pain), 1 (intensity 1), 2 (intensity 2), 3 (intensity 3), 4 (intensities 4-5), and 5 (intensities 6-15). We further evaluate P-MSDA on additional in-the-wild affective computing datasets. Aff-Wild2 [66], where each video is treated as a subject and 10 diverse target videos are selected, with the rest serving as sources. BAH [67] contains self-recorded videos in semi-uncontrolled settings; we use 10 subjects as targets and the remaining as source domains.
@@
IV-B Implementation Detail
In all experiments, we employ the ResNet18 backbone [71], which consists of the encoder $F$ and the discriminative component $C$. To adapt $F$ for subject-based MSDA, we follow the same protocol as  [18] to remove the first ReLU, the MaxPool layers, and the final 2D adaptive average pooling layer. The backbone and classifier are shared across domains, with images resized to $\times100$. Models are trained using SGD with a batch size of 16 and a learning rate of $10^{-4}$. Closest source selection uses $\gamma=0.8$ and $top_{s}=40$ through empirical evaluation; details are provided in Section III-B of the supplementary material. Target pseudo-labels are generated with $\tau=\tau_{0}-\delta\lfloor e/N\rfloor$, where $\tau_{0}=0.90$, $N=20$, and $\delta=0.01$. For ACPL, we follow [18] using horizontal flips as augmentation. Replay memory is updated after each source with $top_{n}$ and $N_{r}$ set to 2000 samples
IV-C Baseline Methods
To evaluate the performance of our method, we define subjects as source and target domains. The first experiment was conducted on the BioVid dataset, where 77 subjects were treated as sources and adapted to the remaining ten target subjects. The following experiment is on the UNBC-McMaster dataset, which includes 20 subjects in the source domain adapted to the remaining five subjects in the target domain. To evaluate the efficacy of our model, we further experimented with a cross-dataset with 20 UNBC-McMaster sources adapted to 10 BioVid target subjects. We follow the previous work of Zeeshan et al. [18] to define the MSDA standard for pain recognition. Source-combined: We first define the lower-bound, which is the traditional approach of training a model by using all the sources and testing the target subject. This approach is also known as source-only, as it does not adapt to the target data. The second experiment combines all subjects as before and then adapts to a target subject as in standard UDA. We compare our method with two standard UDA methods, Domain Adversarial Neural Networks (DANN) [60], Conditional Adversarial Domain Adaptation (CDAN) [61], and one Subject-based [18] UDA method. Multi-source DA: We treat each subject as a separate domain while adapting to the target. We compare our method with three standard and one subject-based MSDA approaches: moment matching for multi-source domain adaptation (M3SDA) [35], implicit alignment (SImpAl) [37], contrastive multi-source domain adaptation (CMSDA), and subject-based domain adaptation [18]. M3SDA technique was the baseline method for MSDA classification tasks that reduced the discrepancy based on the moment-matching approach between domains. CMSDA and SImpAI methods are based on generating the target PLs. Subject-based DA is the STA technique used in multi-source domain adaptation for pain estimation.
V Results and Discussion
V-A Comparison with State-of-the-Art
The result on the BioVid dataset is shown in Table I. In the source-only method, we follow the lower bound approach without any form of domain adaptation, where a model is trained on training subjects (sources) and evaluated on unlabeled target subjects. In the source-combined (blending) setting, we compare our results with subject-based UDA, DANN, and CDAN. Among these, DANN outperforms the other source-combined methods, achieving an average accuracy of 0.65. We compare our technique with four state-of-the-art MSDA methods. SImpAl, CMSDA, M3SDA, and subject-based MSDA. Our method achieves higher performance for all target subjects with an average accuracy of 0.88 that exceeds the baseline by a large margin. The closest result was with a sub-basedtop-k with an average of 0.83 with a similar performance in Sub-2 with 0.86 accuracy. Almost every model performance is impressive in subject Sub-1. However, our method yields remarkable performance in achieving the accuracy of 0.99, which is the same as Oracle. The result on UNBC-McMaster is presented in Table II. In source-combined DANN achieves the highest performance of 0.83, outperforming other baseline methods. In contrast, our approach outperforms all methods on every target subject adaptation, including source-only and state-of-the-art MSDA methods, achieving an average accuracy of 0.88, which is a gain from the previous subject-based method 0.86. In particular, there is a significant improvement in Sub-1 performance with an increase of 0.6 compared to the subject-based approach. For Sub-3, we match the performance with the subject-based method with 0.94 accuracy. Fig. 3 reports the average accuracy across 10 subjects from the Aff-Wild2 and the BAH datasets. Our proposed P-MSDA achieves the best performance, with 0.46 on Aff-Wild2 and 0.71 on BAH, significantly improving over both source-only and subject-based baselines. These gains highlight the robustness of our approach in handling complex affective states across unconstrained, real-world recordings. A more detailed subject-wise breakdown and additional analysis are provided in the supplementary material.
Figure 3: Average accuracy on Aff-Wild2 and BAH datasets.
@@ V-A Comparison with State-of-the-Art
SImpAI [37]
CMSDA [20]
Sub-based [18]
Sub-basedtop-k [18]
@@ V-B Cross-Dataset Evaluation
Figure 4: Comparison of replay sample selection strategies: No Preserve (only new subject), Preserve Random (fixed random samples), Closest k-means (cluster-based), Closest DBSCAN (100 samples/subject), and P-MSDA (Ours) (density-based pertinent samples).
Figure 5: T-SNE visualization of Biovid embeddings across source, replay, and target domains (Subject-1 and Subject-5). The replay domain preserves samples over training steps: the Initial Step retains more source data, while later steps select fewer but more pertinent samples, reducing distant-subject influence and enhancing target alignment.
@@
Figure 7: (a) We compare four techniques to gradually introduce source data; i) Random Source Samples, ii) Closest Source Samples, iii) Closest Source Subjects (all), and iv) Closest Source subjects (ours). Every point represents an average of all the target subjects in BioVid. On the x-axis, each increment is equivalent to approximately 2000 samples or a single subject. (b) Loss when introducing a new source domain, adapting to BioVid target subject-10. (c) Time complexity comparison for training; All Previous Subjects, CMSDA, SImpAI, and P-MSDA (ours).
V-C Impact of Curriculum-Guided Progressive Adaptation
To evaluate the role of introducing sources in a curriculum-guided manner within our progressive adaptation method, we conducted an ablation comparing ordered adaptation (sorted by similarity) against a non-ordered alternative (shuffled adaptation). Specifically, we compared four strategies: (1) Random Source Samples — subsets of $2000$ images randomly selected from all source domains, introduced progressively to simulate a shuffled or non-ordered adaptation; (2) Closest Source Samples — subsets of $2000$ samples selected by cosine similarity from all sources; (3) Closest Source Subjects (All) — introducing multiple similar subjects together and retaining all previously seen subjects; and (4) Closest Source Subjects (Ours) — progressively introducing entire subjects ranked by similarity to the target while retaining only the most relevant past samples via the replay dictionary. As shown in Fig. 6(a), the Random Source Samples baseline yields the weakest performance, confirming that the absence of ordering is suboptimal. Selecting the Closest Source Samples improves over random selection, but still falls short of subject-level ordering, indicating that sample-level similarity alone is insufficient to capture consistent identity-specific cues. While Closest Source Subjects (All) initially performs well, the accumulation of too many past subjects reduces focus on newly introduced sources, leading to diminishing gains. In contrast, our curriculum-based Closest Source Subjects approach achieves the best and most stable improvement over time, demonstrating that adaptation order is critical. By preserving subject-specific contextual information and gradually increasing domain diversity, this strategy allows the model to learn more coherent and relevant representations, leading to faster convergence and better final accuracy.